OpenStudy (anonymous):

limit as x -->-1 from the right (x^2 - 2x +1)/(x+1)

OpenStudy (raden):

i u subtitute x=-1, what happend with that fraction ?

OpenStudy (anonymous):

it becomes 0/0

OpenStudy (raden):

nopes it should becomes ((-1)^2 - 2(-1) + 1) /(-1+1) = (1+2+1)/0 = 4/0

OpenStudy (raden):

so, the value limit it will becomes infinity

OpenStudy (anonymous):

oh silly me, i gotcha thanks a bundle! :3

OpenStudy (raden):

yw